OpenStudy (anonymous):

(A reward will be given to the user with the best response) Prove these identities: (i) Sin (x)/Sec (x) = 1/ Tan (x) + Cos (x) (ii) 1 - Cos (x)/ Sin (x) = Sin (x) / 1 + Cos (x)

OpenStudy (anonymous):

I solved the 2nd one for you if you take them as fractions like this \[\frac{ 1-cosx }{ sinx }=\frac{ sinx }{ 1+cosx }\] Cross multiply and you get \[\sin^2x=1-\cos^2x\] You will see something interesting if you add the cos part \[\sin^2x + \cos^2x = 1\] This is one of the Pythagorean identities meaning both sides are identical to one another